What Z-scores correspond to separating middle 60% of the normal distribution from the rest of the distribution?
Answer #1

middle 60% of score is

P(-z < Z < z) = 0.60

P(Z < z) - P(Z < -z) = 0.60

2 P(Z < z) - 1 = 0.60

2 P(Z < z) = 1 + 0.60 = 1.60

P(Z < z) =1.60 / 2 = 0.8

P(Z <0.84 ) = 0.8

z  ± 0.84 ( using standard normal table)